OpenStudy (anonymous):

Give the value of 'a' in terms of 'b' and 'c' from the following expression: 2a-8b = 12c

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what it wants you to do is solve for 'a'

jimthompson5910 (jim_thompson5910):

have you gotten started at all?

OpenStudy (anonymous):

@jim_thompson5910 sorry no im not sure how to

jimthompson5910 (jim_thompson5910):

we have 2a-8b = 12c

jimthompson5910 (jim_thompson5910):

we want 'a' on its own side all by itself let's have 'a' on the left side

jimthompson5910 (jim_thompson5910):

that must mean we have to move over that 8b term

jimthompson5910 (jim_thompson5910):

how do we do this?

OpenStudy (anonymous):

so itll become 12c-8b all divided by 2 which would then equal 6c-4b=a

OpenStudy (anonymous):

do we then plug that in for a? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

why are you subtracting 8b from both sides?

OpenStudy (anonymous):

oh its adding 8b sorry

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

and yes, after you add 8b to both sides, you divide both sides by 2

OpenStudy (anonymous):

making it all 6c+4b=a

jimthompson5910 (jim_thompson5910):

correct, so a = 4b+6c

OpenStudy (anonymous):

okay, great. Is that all?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome